# Analyzing Fine-tuning Representation Shift for Multimodal LLMs Steering

### v1

@@ Abstract
Multimodal LLMs have reached remarkable levels of proficiency in understanding multimodal inputs, driving extensive research to develop increasingly powerful models. However, much less attention has been paid to understanding and explaining the underlying mechanisms of these models. Most existing explainability research examines these models only in their final states, overlooking the dynamic representational shifts that occur during training. In this work, we systematically analyze the evolution of hidden state representations to reveal how fine-tuning alters the internal structure of a model to specialize in new multimodal tasks. Using a concept-based approach, we map hidden states to interpretable visual and textual concepts, enabling us to trace changes in encoded concepts across modalities as training progresses. We also demonstrate the use of shift vectors to capture these concepts changes. These shift vectors allow us to recover fine-tuned concepts by shifting those in the original model. Finally, we explore the practical impact of our findings on model steering, showing that we can adjust multimodal LLMs behaviors without any training, such as modifying answer types, captions style, or biasing the model toward specific responses. Our work sheds light on how multimodal representations evolve through fine-tuning and offers a new perspective for interpreting model adaptation in multimodal tasks. The code for this project is publicly available at https://github.com/mshukor/xl-vlms.
@@ 1 Introduction
Figure 1: Overview. We analyse the concepts representations change due to fine-tuning. This has direct implications on steering MLLMs without any extra training.
@@
With the rapid progress in Large Language Models (LLMs) [7, 10, 34, 52, 22], Multimodal LLMs (MLLMs) [11, 30, 27, 3] have recently demonstrated remarkable capabilities in addressing complex multimodal tasks such as image captioning and visual question-answering.
@@
MLLMs are typically composed of a visual encoder, an LLM, and an intermediary connector. Following initial unimodal pretraining—and, in many cases, multimodal pretraining on large datasets—these models can be further specialized by training on multimodal datasets [26, 1]. Given the high computational cost of training these models, recent research has proposed more efficient approaches, such as creating diverse, high-quality instruction-tuning datasets [30] or keeping the LLM frozen and fine-tuning small amount of model parameters, like the connector [47, 31, 54, 46]. These approaches take advantage of the impressive ability of frozen LLMs to generalize to multimodal data [45]. Despite the different efficient tuning methods, training these models still incurs significant cost due to backpropagating the gradient through large models.
@@
While substantial progress has been made in developing high-performing MLLMs, relatively few studies aim to understand them [37, 44, 48, 4, 60, 45]. Existing work typically conducts post-hoc analyses of MLLMs in isolation, overlooking the internal changes during fine-tuning. Research by [45] addresses this gap to some extent by examining the internal multimodal alignment as it evolves during training.
In contrast, we focus on exploring how MLLMs update their internal semantic representations when fine-tuned for multimodal tasks (see Fig. 1). Understanding how fine-tuning alters learned concepts—and, consequently, the features encoded by the model—opens the possibility of directly modifying these features to steer model outputs, reducing the need for additional training and its associated costs.
@@
Our approach leverages concept-based explainability to investigate how concept representations evolve from the original model to its fine-tuned version. We find that fine-tuning on a specific task reshapes these concepts, with some adjusting subtly to align with the task, and others emerging or disappearing altogether (see Fig. 2). Notably, most fine-tuned concepts can be reconstructed from the original model by translating its original concepts along specific shift vectors. Lastly, we explore the implications of this study for steering MLLMs, demonstrating how model responses can be modified without additional training. Our key findings are summarized as follows:
  • We show that learned concepts adapt specifically to the fine-tuning task, with some concepts becoming more specialized and others undergoing complete transformation.
  • Many of the concepts encoded in the fine-tuned model can be reconstructed from those in the original model using shift vectors in the latent space.
  • We apply these findings to steer MLLMs (e.g., LLaVA) outputs, demonstrating the possibility to steer the model’s answers or response style. To the best of our knowledge, this is the first work exploring MLLMs steering.
@@ 2 Related Work
Concept-based explainability. Concept-based explainability methods have emerged as an alternative to traditional feature attribution based methods, that are capable of extracting key semantic features from the model internal representations. Most post-hoc concept explainability approaches are based on the idea of concept activation vectors (CAV) [23], which represent concepts as vectors in the activation space. Instead or relying on human annotations, recent works have proposed methods to automatically discover concepts via clustering [17, 57] or matrix decomposition [14], which can be viewed as instances of a dictionary learning problem [13]. Initially focusing on understanding vision models, dictionary learning for concept extraction has been extended to LLMs, for instance, in the form of sparse autoencoders [21, 41]. However, none of the prior approaches have been applied to understand multimodal models, with the exception of recently proposed CoX-LMM [37].
@@
MLLMs and Explainability. Multimodal LLMs [30, 3, 27, 54] have recently garnered significant interest. These models generally adopt a late fusion architecture, which includes an image encoder [40, 58, 15], a connector module, and an LLM [52, 22, 50]. This family of models has inspired extensive research to better understand them and explain their behavior. For example, studies like [44, 35, 20] seek to identify multimodal neurons within LLMs or analyze modality-specific sub networks [45]. Given that these models are text-generative, some methods leverage this property to simply generate textual explanations for model outputs [48, 56, 16, 9]. Based on LLMs, multimodal models benefit from in-context learning capabilities, which have been examined for limitations, including biases [4] and links to hallucinations [48], as well as the factors that may enhance their in-context learning performance [8, 39]. Related to our approach, CoX-LMMs [37] employs dictionary learning to extract multimodal semantic concepts from model representations. However, these studies typically assess models only in their final trained states, overlooking the dynamic changes that occur during training. Only limited works, such as [45], have investigated this training process, focusing specifically on implicit alignment between image and text modalities. In this work, we investigate how multimodal concepts within the model evolve throughout fine-tuning and explore the implications of these shifts on model steering.
@@
Steering models with feature editing. In contrast to editing model weights, representation or feature editing methods [55, 49, 53] aim to modify model outputs without altering the model’s weights. A prominent approach within this family involves identifying steering vectors, or directions in the feature space (often within the residual stream), that are linked to contrasting concepts. These methods have been applied to language models for various purposes, such as enhancing factuality or reducing hallucinations [36], inducing sentiment shifts or detoxification [53, 51], improving refusals to harmful requests [2], promoting truthfulness by modifying the output of attention heads [28], and erasing specific concepts or biases [6, 42]. However, these techniques have primarily focused on language models, with their application to MLLMs yet to be explored.
3 Notations and Background
This section covers the background and the notations.
Model architecture.
@@ 3 Notations and Background
A generic MLLM consists of a visual encoder $f_{V}$, a trainable connector $C$, and a language model $f_{\text{LM}}$. We assume that the model is pretrained on a multimodal dataset $\mathcal{S}=\{(x_{i},y_{i})\}_{i}$, such as captioning, where $x_{i}\in\mathcal{X}$ represents images and $y_{i}\subset\mathcal{Y}$ are the associated captions specified as sequence of tokens from token vocabulary space $\mathcal{Y}$. The model is trained to generate the next text tokens, conditioned on text and images. The input to $f_{\text{LM}}$ is a sequence of tokens that includes the concatenation of: (1) $N_{V}$ visual tokens extracted from the image $x$ via the visual encoder followed by the connector ($C(f_{V}(x))$), and (2) linearly embedded textual tokens corresponding to the text instruction and previously predicted tokens. This can be expressed as:
@@
Figure 2: Analysis framework overview. We analyse the concepts representation change during fine-tuning and derive shift vectors that can recover most of the concept in the fine-tuned model.
Concept-based explainability.
To understand the internal representations of any given MLLM $f$, we leverage the approach introduced in [37]. Specifically, for a token of interest TOI, we first extract the residual stream representations, after the $l-$th layer, $h_{(l)}^{p}={f_{l}}(x)$ [12] of a subset of samples $X_{\textit{TOI}}\subset\mathcal{S}$. $X_{\textit{TOI}}$ includes all the samples $x$ that contain TOI in both the ground truth and predicted response ($\textit{TOI}\in\hat{y};\hat{y}=f(x)$). For simplicity, given a sample $x_{m}$ the extracted hidden states (or features) corresponding to the TOI token is denoted as $\bm{z}_{m}={f_{l}}(x_{m})\in\mathbb{R}^{D}$, where $D$ is the features dimension. Thus the features from all samples gives a feature matrix $\bm{Z}\in\mathbb{R}^{D\times M}$ where $M=|X_{\textit{TOI}}|$. The feature matrix $\bm{Z}$ is then decomposed as $\bm{Z}\approx\bm{UV}$ to discover the encoded concepts. Here, $\bm{U}\in\mathbb{R}^{D\times K}$ is the matrix of $K$ concepts and $\bm{V}\in\mathbb{R}^{K\times M}$ represents the coefficients/activations of the samples projected onto these concepts. Each column $\bm{u}_{k}\in\bm{U}$ corresponds to a concept, while each column of $\bm{V}$ encodes the activation of these concepts for a given sample. Note that any given representation $f_{l}(x)$ can be projected on $\bm{U}$ to obtain its activation vector $\bm{v}(x)\in\mathbb{R}^{K}$, i.e. $f_{l}(x)\approx\bm{U}\bm{v}(x)$.
Each extracted concept is interpreted through grounding in both image and text spaces. Specifically, the top $N_{\text{MAS}}$ most activating samples for the concept $\bm{u}_{k}$ represent its image grounding $\bm{X}_{\text{MAS}}$:
@@
We employ $K$-Means to learn our concept dictionaries. This is motivated by $K$-Means’ simplicity, and straightforward arithmetic manipulation of the clusters/concepts it allows, making it a more suitable for our analysis.
Fine-tuning setup.
To develop a specialized model $f^{b}$ tailored for a particular task—or, specifically, for a set of target concepts—the original model $f^{a}$ is fine-tuned on samples that include a set of words, $\{w_{1},\cdots,w_{m}\}$, associated with these target concepts. Our focus is on fine-tuning the LLM component, using Low-Rank Adaptation (LoRA) [19], an approach known to deliver performance comparable to full model fine-tuning while being significantly more efficient [30, 27]. After finetuning, we extract feature matrices $\bm{A}\in\mathbb{R}^{D\times M^{a}}$ ($M^{a}=|X_{\textit{TOI}}^{a}|$) from $f^{a}$ and $\bm{B}\in\mathbb{R}^{D\times M^{b}}$ ($M^{b}=|X_{\textit{TOI}}^{b}|$) from $f^{b}$, which contain the set of residual stream representations composed of:
$\bm{a_{i}}=f_{l}^{a}(x^{a}_{i}),\quad\bm{b_{j}}=f_{l}^{b}(x^{b}_{j}),$
where $x^{a}_{i}\in X_{\textit{TOI}}^{a}$ and $x^{b}_{j}\in X_{\textit{TOI}}^{b}$ are samples where TOI is present in the ground truth and the response of both $f^{a}$ and $f^{b}$. Next, the matrices $\bm{A}$ and $\bm{B}$ are decomposed to extract the concepts each encodes as:
$A\approx\bm{U}^{a}\bm{V}^{a},\quad B\approx\bm{U}^{b}\bm{V}^{b},$
where the hidden states of each sample can be represented as $\bm{a}_{i}=\bm{U}^{a}\bm{v}^{a}(x^{a}_{i})$ and $\bm{b}_{j}=\bm{U}^{b}\bm{v}^{b}(x^{b}_{j})$, $\bm{U}^{a}$ with $\bm{U}^{b}$ $\in\mathbb{R}^{D\times K}$ representing the learned concepts $\{u^{a}_{k}\}_{k=1}^{K},\{u^{b}_{k}\}_{k=1}^{K}$ respectively , and $\bm{V}^{a}$ and $\bm{V}^{b}$ $\in\mathbb{R}^{K\times M}$ are the corresponding activations. Finally, we ground the concepts in the image modality (Eq. 1) to obtain $\bm{X}_{k,\text{MAS}}^{a}$ and $\bm{X}_{k,\text{MAS}}^{b}$ as well as in the text modality (Eq. 2) to obtain $\bm{T}_{\text{words}}^{a}$ and $\bm{T}_{\text{words}}^{b}$.
Concepts similarity.
To quantify the similarity of two concepts $u$ and $u^{\prime}$, we define Text Grounding Overlap between them as:
@@
One can similarly define the Image Grounding Overlap and discussed in the Appendix.
4 Fine-tuning and evolution of concept representations
We aim to study how fine-tuning process affects the concepts learned by the model. Then we try to recover the fine-tuned model concepts using shift vectors computed in the feature space.
[…]
Implementation details.
We focus on MLLMs with the architecture described in Section 3. The image encoder $f_{V}$ is a ViT-L/14 CLIP [40], followed by a transformer-based connector that reduce the number of encoded visual tokens. The LLM $f_{\text{LM}}$ is an OPT-6.7B model [59] with 32 layers. We conduct our study in a controlled setups that consists of specializing the model on a target dataset. Specifically, we fine-tune on three different subsets of the Visual Genome dataset [24], related to places, colors, and sentiments (see App. A for more details).
4.1 Impact of fine-tuning on learned concepts
The fine-tuning process introduces changes in the overall structure of the learned concepts. This change can be observed through the concepts groundings in both the text and image space (Fig. 3).
[…]
Matched concepts.
To analyze how each concept changes after fine-tuning, we focus on each concept and its match (see Fig. 4). Specifically, we define a matching function $m:i\rightarrow j^{*}$ which associates each concept vector $\bm{u}^{a}_{i}$ in set $\bm{U}^{a}$ to its closest vector $\bm{u}^{b}_{j^{*}}$ in set $\bm{U}^{b}$ based on cosine similarity, i.e:
$m(i)=\operatorname*{arg\,max}_{\bm{u}^{b}_{j}\in\bm{U}_{b}}\text{cos}(\bm{u}^{a}_{i},\bm{u}^{b}_{j}),\;\text{cos}(\bm{u}^{a}_{i},\bm{u}^{b}_{j})=\frac{\bm{u}^{a}_{i}\cdot\bm{u}^{b}_{j}}{\|\bm{u}^{a}_{i}\|\|\bm{u}^{b}_{j}\|}$ (4)
Figure 5: Text grounding overlap between original and fine-tuned models. Illustration of T-Overlap between original and matched fine-tuned model concepts.
Concept evolution.
[…]
Figure 6: Concepts evolve differently due to fine-tuning. Left: concepts extracted from the original model $f^{a}$. Right: concept extracted from $f^{b}$, fine-tuned to focus on places. Each concept is grounded in image and text.
We also notice that the T-Overlap decreases with the number of training iterations, indicating that fine-tuning leads to deviation from original model concepts (see App. A for more details).
4.2 Recovering concepts in fine-tuned model via shift vectors
In this section, we study if the learned concepts in the fine-tuned model can be recovered from the original model concepts, without any training or decomposition. In particular, we propose to characterize the change in original concepts due to fine-tuning as linear directions in the feature space, referred to these as concept shift vectors.
Shifting original model concepts.
To compute the concept shift vectors, we first associate each concept $\bm{u}^{a}_{k}$ in the original model with a subset of samples common to the analysis of both $f^{a}$ and $f^{b}$. Specifically, projecting the representation of each sample $x_{m}\in\bm{X}_{TOI}^{a}\cap\bm{X}_{TOI}^{b}$ on $\bm{U}^{a}$ and selecting those that activate $\bm{u}^{a}_{k}$ the most:
@@ 4.2 Recovering concepts in fine-tuned model via shift vectors
where $\alpha$ is a coefficient to control the shift magnitude. Unless otherwise stated, we use $\alpha=1$ as the default magnitude of shift. It is worth noting that given the concept shift vectors, the computation of shifted concepts does not rely on accessing the fine-tuned model.
Evaluating fine-tuned concept recovery.

### v2

@@ Abstract
Multimodal LLMs (MLLMs) have reached remarkable levels of proficiency in understanding multimodal inputs. However, understanding and interpreting the behavior of such complex models is a challenging task, not to mention the dynamic shifts that may occur during fine-tuning, or due to covariate shift between datasets. In this work, we apply concept-level analysis towards MLLM understanding. More specifically, we propose to map hidden states to interpretable visual and textual concepts. This enables us to more efficiently compare certain semantic dynamics, such as the shift from an original and fine-tuned model, revealing concept alteration and potential biases that may occur during fine-tuning. We also demonstrate the use of shift vectors to capture these concepts changes. These shift vectors allow us to recover fine-tuned concepts by applying simple, computationally inexpensive additive concept shifts in the original model. Finally, our findings also have direct applications for MLLM steering, which can be used for model debiasing as well as enforcing safety in MLLM output. All in all, we propose a novel, training-free, ready-to-use framework for MLLM behavior interpretability and control. Our implementation is publicly available.  [^1] Project page and code: https://pegah-kh.github.io/projects/lmm-finetuning-analysis-and-steering/
$\star$\$$\star$\$footnotetext: First authors
@@ 1 Introduction
With the rapid progress in Large Language Models (LLMs) [7, 10, 44, 65, 28], Multimodal LLMs (MLLMs) [12, 39, 34, 3, 61] have recently demonstrated remarkable capabilities in addressing complex multimodal tasks such as image captioning and visual question-answering.
@@
MLLMs are typically composed of a visual encoder, an LLM, and a connector. Following initial unimodal pretraining—and, in many cases, multimodal pretraining on large datasets—these models can be further specialized by training on multimodal datasets [33, 1]. Given the high computational cost of training these models, recent research has proposed more efficient approaches, such as creating diverse, high-quality instruction-tuning datasets [39] or keeping the LLM frozen and fine-tuning small amounts of parameters, like the connector [58, 41, 67, 57]. These approaches take advantage of the ability of frozen LLMs to generalize to multimodal data [56]. Despite the different efficient tuning methods, training these models still incurs significant costs.
@@
While substantial progress has been made in developing high-performing MLLMs, relatively few studies aim to understand them [47, 55, 59, 4, 73, 56, 61]. Existing work typically conducts post-hoc analyses of MLLMs in isolation, overlooking the internal changes due to fine-tuning. Research by [56] addresses this gap to some extent by examining the internal multimodal alignment as it evolves during training.
In this work, we apply concept-level analysis to provide a readable understanding of MLLM behavior and, in particular, semantic dynamics that may occur due to fine-tuning, or due to covariate shift when considering different datasets.
@@
In the first case, we find that fine-tuning on a specific task potentially reshapes learned latent concepts, with some adjusting subtly to align with the task, and others emerging or disappearing altogether (see Fig. 1). Notably, we find that most fine-tuned concepts can be reconstructed from the original model by translating its original concepts in the direction of specific concept shift vectors, reducing the need for additional training and its associated costs. Furthermore, we explore the implications of the proposed analysis for MLLMs steering, demonstrating how model outputs can be modified inexpensively without additional training. Our key findings are summarized as follows:
  • We apply latent concept-level analysis to provide readable understanding on MLLMs’ behavior ; in particular, we show that fine-tuning can introduce significant alteration in the original concepts.
  • We show that we can control the MLLM’s behavior w.r.t. certain concepts by simply manipulating shift vectors.
  • Lastly, our findings also have direct applications for steering MLLM outputs, which find use for model debiasing as well as safety control.
In a nutshell, we propose a novel, ready-to-use framework (including code) for MLLM behavior interpretability and control, debiasing and steering, which, we believe, will pave the way for future research.
Figure 1: Framework overview. We apply concept-level analysis for MLLM behavior monitor and control, for (left) understanding and manipulating (through shift vectors) concept changes due to fine-tuning, as well as (right) MLLM steering for debiasing or safety control.
@@ 2 Related Work
Concept-based explainability. Concept-based explainability methods have emerged as an alternative to traditional feature attribution based methods, that are capable of extracting key semantic features from the model internal representations. Most post-hoc concept-based approaches are based on the idea of concept activation vectors (CAV) [30], which represent concepts as vectors in the activation space. Instead or relying on human annotations, recent works have proposed methods to automatically discover concepts via clustering [21, 71] or matrix decomposition [18], which can be viewed as instances of a dictionary learning problem [17]. Initially focusing on understanding vision models, dictionary learning for concept extraction has been extended to LLMs e.g. using sparse autoencoders [27, 52]. However, none of the prior approaches have been applied to understand MLLMs, with the exception of recently proposed CoX-LMM [47].
@@
MLLMs and Explainability. Multimodal LLMs [39, 3, 34, 67] have recently garnered significant interest. They typically adopt a late fusion architecture, and consist of an image encoder [51, 72, 19], a connector, and an LLM [65, 28, 63]. This family of models has inspired extensive research to better understand them and explain their behavior. For example, studies like [55, 45, 26] seek to identify multimodal neurons within LLMs or analyze modality-specific sub networks [56]. Some methods leverage the fact that these models are text-generative to simply generate textual explanations for model outputs [59, 70, 20, 8]. MLLMs benefit from in-context learning capabilities, which have been examined for limitations, including biases [4] and links to hallucinations [59], as well as the factors that may enhance their in-context learning performance [9, 49]. Related to our approach, CoX-LMMs [47] employs dictionary learning to extract multimodal semantic concepts from model representations. However, these studies typically assess models only in their final trained states, overlooking the dynamic changes that occur during training. Only limited works, such as [56], have investigated explaining changes due to fine-tuning, focusing specifically on implicit alignment between image and text modalities. In this work, we investigate how multimodal concepts within the model evolve throughout fine-tuning and explore the implications of these shifts on model steering.
@@
Steering models with feature editing. In contrast to editing model weights, representation or feature editing methods [69, 62, 66] aim to modify model outputs without altering the model’s weights. A prominent approach within this family involves identifying steering vectors, or directions in the feature space (often within the residual stream), that are linked to contrasting concepts. These methods have been applied to language models for various purposes, such as enhancing factuality or reducing hallucinations [46], inducing sentiment shifts or detoxification [66, 64], improving refusals to harmful requests [2], promoting truthfulness by modifying the output of attention heads [35], and erasing specific concepts or biases [6, 53]. However, their application to MLLMs is yet to be explored. Another set of approaches related to steering methods are based on In-context learning (ICL) [15, 16, 29], where prompts are carefully designed to induce desired behavior. Yet, ICL requires predefined demonstrations and lacks interpretability at a concept level. In contrast, our method offers lightweight steering, extending such capabilities to MLLMs without requiring any training, for instance as in ReFT [69].
3 Methodology Overview
Our framework is summarized in Fig. 1. We apply concept-level analysis of MLLM latent space (Section 3.1) to define and compare concepts between different setups. This allows us to monitor conceptual changes occurring during fine-tuning and manipulate MLLM behavior at a conceptual level (Section 3.2). Lastly, our framework finds applications for MLLM steering for e.g. debiasing and safety control (3.3) with negligible computational burden.
@@ 3 Methodology Overview
MLLM setup. A generic MLLM consists of a visual encoder $f_{V}$, a trainable connector $C$, and a language model $f_{\text{LM}}$. We assume that the model is pretrained on a multimodal (e.g. captioning) dataset $\mathcal{S}=\{(x_{i},y_{i})\}_{i}$, where $x_{i}\in\mathcal{X}$ represents images and $y_{i}\subset\mathcal{Y}$ are the associated captions specified as sequence of tokens from token vocabulary space $\mathcal{Y}$. The model is trained to generate the next text tokens, conditioned on text and images. The input to $f_{\text{LM}}$ is a sequence of tokens that includes the concatenation of: (1) $N_{V}$ visual tokens extracted from the image $x$ via the visual encoder and connector ($C(f_{V}(x))$), and (2) linearly embedded textual tokens corresponding to the text instruction and previously predicted tokens. This can be expressed as:
@@
3.1 Retrieval and comparison of latent concepts
To understand the internal representations of any given MLLM $f$, we leverage the approach introduced in [47]. Specifically, given a set of $M$ images $\{x_{1},...,x_{M}\}$ we extract a set of residual stream representations from some layer $l$ of the MLLM $f$. These representations $\bm{z}_{m}={f_{l}}(x_{m})\in\mathbb{R}^{D}$ (one per image) are collected in a feature matrix $\bm{Z}\in\mathbb{R}^{D\times M}$. Typically, the set of images and extracted representations correspond to a particular token of interest $TOI$ (e.g., ‘Dog’, ‘Cat’, ‘Person’, etc.) in the predicted caption. However, the extraction can also be performed for a larger set of target tokens. This feature matrix $\bm{Z}$ is then decomposed as $\bm{Z}\approx\bm{UV}$ to recover the concepts in the latent embedding space. Here, $\bm{U}\in\mathbb{R}^{D\times K}$ is the matrix of $K$ concepts and $\bm{V}\in\mathbb{R}^{K\times M}$ represents the coefficients/activations of the samples projected onto these concepts. Different decompositions of the matrix $K$ result in various concepts, inheriting the properties of the decomposition (such as low grounding overlap with PCA). We employ $K$-Means to learn our concept dictionaries. This is motivated by $K$-Means’ simplicity, and straightforward arithmetic manipulation of the clusters/concepts it allows. Each column $\bm{u}_{k}\in\bm{U}$ corresponds to a concept, while each column of $\bm{V}$ encodes the activation of these concepts for a given sample. Note that any given representation $f_{l}(x)$ can be projected on $\bm{U}$ to obtain its activation vector $\bm{v}(x)\in\mathbb{R}^{K}$, i.e. $f_{l}(x)\approx\bm{U}\bm{v}(x)$. Each extracted concept is then interpreted through grounding in both image and text spaces. Specifically, the top $N_{\text{MAS}}$ that activates concept $\bm{u}_{k}$ the most represent its image grounding:
@@ 3.1 Retrieval and comparison of latent concepts
Finally, to quantify the similarity of two concepts (i.e. the columns of $\bm{U}$ we define the Text Grounding Overlap as:
@@
Now that we have defined a generic framework, we present two subcases for extracting and manipulating concepts.
3.2 Evolution of concepts through fine-tuning.
Setup overview. An original model $f^{a}$ is typically fine-tuned to produce a specialized model $f^{b}$ for a particular task—or, specifically, for a set of target concepts. This fine-tuning can be conducted on samples that include a set of words $\{w_{1},\cdots,w_{m}\}$ associated with these target concepts. For instance, if we fine-tune a image captioning model to emphasize colors in the image, the set of words will simply be these colors. Efficient fine-tuning is typically achieved using Low-Rank Adaptation (LoRA) [25, 39, 34]. Fine-tuning can selectively alter certain representations, leading to shifts in the conceptual space encoded by the model. Using the interpretability framework discussed in Section 3.1 we can study these shifts at a readable conceptual level.
Concept recovery via shift vectors. To study the change from an original model $f^{a}$ to a finetuned model $f^{b}$, we fix the dataset $S^{(1)}=S^{(2)}$, and obtain two sets of embeddings from $f^{a},f^{b}$ respectively, i.e. $A\approx\bm{U}^{a}\bm{V}^{a}$, $B\approx\bm{U}^{b}\bm{V}^{b}$, where $\bm{U}^{a},\bm{U}^{b}\in\mathbb{R}^{D\times K}$ are $K$ concepts extracted from each model. We propose to characterize the concept changes from an original to fine-tuned model as linear directions in embedding space or concept shift vectors. To do so, we first associate each original concept $\bm{u}^{a}_{k}\in\bm{U}^{a}$ with a subset of samples where $\bm{u}^{a}_{k}$ is the most activated concept:
@@
where $\alpha$ is a coefficient to control the shift magnitude. Unless otherwise stated, we use $\alpha=1$ as the default magnitude of shift. It is worth noting that given the concept shift vectors, the computation of shifted concepts does not rely on accessing the fine-tuned model. Practically speaking, this means that we can ”push” the original model towards the concepts of a fine-tuned one with very little overhead, by simply shifting its latent representation in the direction of the shift vector, as it will be illustrated in the experiments.
3.3 Concept evolution across datasets and applications to model steering
Concept comparison between different datasets. Another interesting subcase of the proposed framework consists in evaluating the shift from one dataset $S^{(1)}$ to another $S^{(2)}$, using the same model and encoding $h$. Beyond interpretability of a model behavior, it also find applications for model steering. Model steering (see Fig. 1 (right)) refers to guiding the model outputs towards desired outcomes by modifying the features without altering the model weights.
Coarse-grained model steering. In coarse-grained or global steering, the objective is to adjust the model outputs $\hat{y}$ to generally align with a set of target samples (e.g., changing answers type). Given input-output samples, we first extract the answer representations $\bm{B}={\bm{b}_{1},...,\bm{b}_{N}}$ at layer $l$ from the target set. Similarly, we obtain representations for the original set $\bm{A}={\bm{a}_{1},...,\bm{a}_{M}}$ (e.g. randomly drawn from the train set). We then compute the coarse steering vector $\bm{s}_{c}$ as:
$\bm{s}_{c}=\frac{\sum_{i}^{N}\bm{b}_{i}}{N}-\frac{\sum_{i}^{M}\bm{a}_{i}}{M},$ (5)
$s_{c}$ is applied to all the samples in the validation set. For instance, the activations $f_{l}(x_{i})$ of a sample $x_{i}$ become:
$\tilde{f_{l}}(x_{i})=f_{l}(x_{i})+\alpha\bm{s}_{c}$ (6)
where $\alpha$ controls the steering strength and it is set to 1 (we study $\alpha$ in Section B.5.3). Thus, in this setup, all examples are coarsely steered in the direction of the steering vector $\bm{s}_{c}$ at layer $l$, before passing $\tilde{f_{l}}(x_{i})$ through the rest of layers and $\tilde{f_{l}}(x_{i})$ becomes the input to the next layer $l+1$.
Fine-grained steering. Unlike global steering, fine-grained steering consists in finding and editing directions that adjust only certain concepts to other ones. To do this, we decompose the hidden states of a set of samples into a set of concepts $\bm{U}$ as previously explained. We then compute a set of fine-grained steering vectors $\bm{s}^{f}={\bm{s}_{11}^{f},...,\bm{s}_{NN}^{f}}$, with $\bm{s}_{ij}^{f}=\bm{s}_{ij}^{f}=\bm{u}_{j}-\bm{u}_{i}$ the steering vector from concept $\bm{u}_{i}$ to $\bm{u}_{j}$. However, not all steering vector are meaningful: options for finding the relevant ones include proximity matching, as well as identifying vectors that have the strongest impact on guiding the model towards generating specific answers or concepts (e.g. producing significantly more target answers). This is more detailed in Section B.3. Various applications of MLLM steering can be explored, such as gender debiasing and safety alignment, as it will be shown below.
4 Experiments
4.1 Fine-tuning experiments
In this section, we study how fine-tuning introduces changes in the overall structure of the learned concepts in MLLMs ( with the architecture described in Section 3).
Implementation details. The main paper covers experiments on the popular LLaVA [38] model comprising a CLIP image encoder, a two-layer MLP connector, and a 7B Vicuna-1.5 LLM. More experiments on a different multimodal model can be found in App. A. We conduct our study in a controlled setup that consists in specializing the model on a target dataset. Specifically, we apply fine-tuning on three different subsets of the Visual Genome dataset [31], related to places, colors, and sentiments (more details in Section A.2).
[…]
Impact of fine-tuning on learned concepts. Fig. 2 depicts this concept change. Throught the concepts groundings for an exemple TOI (person) we see that the fine-tuned model puts a stronger emphasis on places, which is expected and serves as a sanity check for our method.
[…]
Matched concepts. To further analyze how each concept changes after fine-tuning, we focus on each concept and its match. Specifically, we define a matching function $m:i\rightarrow j^{*}$ which associates each concept vector $\bm{u}^{a}_{i}$ in set $\bm{U}^{a}$ to its closest vector $\bm{u}^{b}_{j^{*}}$ in set $\bm{U}^{b}$ based on cosine similarity, i.e $m(i)=\operatorname*{arg\,max}_{\bm{u}^{b}_{j}\in\bm{U}_{b}}\text{cos}(\bm{u}^{a}_{i},\bm{u}^{b}_{j})$.
Fig. 3 shows the text groundings for various concepts, displaying the words with a frequency lower than 5 across concepts (e.g., filtering out high-frequency terms like bus, vehicle, etc.). We observe the emergence of place-related terms across the identified elements, while the overall thematic structure remains consistent. Note that certain concepts may converge toward the same fine-tuned concept. We also analyze how the distances between matched concepts evolve during fine-tuning in Section A.3.
Figure 4: Text grounding overlap (T-Overlap) between original and fine-tuned model concepts. Different concepts change to different extents depending on the fine-tuning.
[…]
Figure 5: Concepts evolve differently due to fine-tuning. Left: concepts extracted from the original model $f^{a}$. Right: matched concept extracted from $f^{b}$, fine-tuned to focus on places. Each concept is grounded in image and text. We observe different levels of adaptation across concepts: some concepts specialize further by adding to place-related words (e.g., Top), some concepts introduce place-related words while staying aligned with the original concept in textual or visual groundings (e.g., Middle), while others undergo complete transformation, diverging significantly from their original meaning to fully embrace place-related elements (e.g., Bottom).
We also notice that T-Overlap decreases with the number of training iterations, indicating that fine-tuning leads to deviation from the original concepts (more details in Section A.3).